OpenStudy (anonymous):

which of the following expressions is equivalent to 1/a^-5? A. 1/a^5 B. a^5 C. a^-5 D. -1/a^5

OpenStudy (rulnick):

B.

OpenStudy (rulnick):

1/a^-5 = 1/(1/a^5) = a^5.

OpenStudy (anonymous):

Rulnick is correct.

OpenStudy (rulnick):

Thank you JuanitaM.

OpenStudy (anonymous):

Just giving encouragement.

OpenStudy (anonymous):

B. a^5